OpenStudy (anonymous):

Is the answer: 43/4? Find the slope intercept form of the equation of the line through the point (0,9) perpendicular to the graph of y=2x+1 1/3 (keep in fraction form)

OpenStudy (tkhunny):

The Slope-Intercept Form is an equation. The correct response cannot possibly be "44/3".

OpenStudy (anonymous):

could you type the question again ?

OpenStudy (anonymous):

My nephew brought this over tonight for help. Neither me or my husband can remember how to do this.

OpenStudy (anonymous):

what is the equation of line given ? is it 

OpenStudy (anonymous):

it is unclear whether you typed 

OpenStudy (anonymous):

No, it's y=2X+1and 1 third

OpenStudy (anonymous):

Find the slope-intercept form of the line through the point (0,9), perpendicular to the graph of y=2X+1and one third

OpenStudy (tkhunny):

Slope-Intercept Form looks like this: y = mx + b You need a slope, m, and an intercept, b, in order to solve the problem. The slope is ALMOST given. It's only a hint. Perpendicular to y = 2x + something. This slope is 2. A perpendicular slope needs to be -1/2 so that (-1/2)*2 = -1. That -2 is a little bit magic. Okay, now we know the new line lies like this: y = (-1/2)x + b You need to follow this far before the last piece.

OpenStudy (anonymous):

1 and 1/3 = 4/3 , that should be easy to do based on what tk***y as suggested

OpenStudy (anonymous):

Yes, but I have been out of school for 20+ years, lol!

OpenStudy (tkhunny):

* That "-1" is a little bit magic. Sorry about that typo.

OpenStudy (anonymous):

y=2x+4/3, slope of this line is say m = 2 lets says slope of line perpendicular to it is say k if lines are perpendicular ,product of slopes is = -1 i.e. m*k=-1 so slope of line perpendicular to it is k=-1/2

OpenStudy (anonymous):

lets say equn of line that is perpendicular is y=kx+C we know its slope k = -1/2 and it passes via (0,9)

OpenStudy (anonymous):

9=0+c so, c=9 eqn of this line is y=-1/2 * x +9

OpenStudy (anonymous):

or y=-0.5x+9

OpenStudy (anonymous):

but since you want it in fraction form you will leave it as Y= (-1/2)X+9

OpenStudy (anonymous):

Oh thank you, thank you!

OpenStudy (anonymous):

I have given the method, but verify the answer, typing fast , may have made some errors but the METHOD is what matters

OpenStudy (anonymous):

(0,9) so what are these numbers?

OpenStudy (anonymous):

(0,9) IS THE POINT THROUGH which the perpendicular line PASSES

OpenStudy (anonymous):

CAPS have been used to highlight the points, no offense intended

OpenStudy (tkhunny):

They are directions for finding your way around a 2-dimensional coordiante system. Typically, the x value is lined up on a horizontal axis (positive to the right) and the y values is lined up on a vertical axis (positive going up). (0,9) means 0 in the x direction and 9 in the positive y direction.

OpenStudy (anonymous):

But do I remember correctly labeling the top of those numbers with alphabet letters? You know, in order to use them?

OpenStudy (anonymous):

So it's x and y

OpenStudy (tkhunny):

That's it. (x,y)

OpenStudy (anonymous):

It's slowly coming back to me. Thank you for your help!

OpenStudy (tkhunny):

Think about it as you go to sleep. It will feel better in the morning. Okay, stop laughing!!

OpenStudy (anonymous):

Ugh, his assignment isn't due until Friday so we can look at his other problem tomorrow.

OpenStudy (tkhunny):

There you have it, then. Time to dream of coordinate axes!!

OpenStudy (anonymous):

I paid my dues in high school and college! I'm dreaming of shopping these days!

OpenStudy (tkhunny):

(sigh) Gone over to the dark side, you have!

OpenStudy (anonymous):

Thank you for putting up with me!